# Ontology Neural Network and ORTSF: A Framework for Topological Reasoning and Delay-Robust Control

### v2

@@ Conclusion:
$d_{\operatorname{PH}}(G_{\mathcal{C}}(t),G_{\mathcal{C}}(t+\delta))\leq L_{c}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+C_{sem}\sqrt{\mathcal{L}_{\operatorname{ph}}}$ (160)
@@
B-B Patch-B: Discrete Predictor Continuity and Grönwall Bound (Complete Proof)
@@
$\|\mathcal{R}_{\operatorname{trace}}(t)-\mathcal{R}_{\operatorname{trace}}(t-\Delta t)\|\leq C_{PH}\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+\eta(\mathcal{L}_{\operatorname{ph}})$ (174)
@@
$\leq L_{\operatorname{total}}(C_{PH}\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+\eta(\mathcal{L}_{\operatorname{ph}}))$ (176)
@@
B-C Patch-D: Robust Phase Margin with Frequency Drift (Complete Proof)
@@
$\frac{d}{df}\log|L(j2\pi f)|\Big{|}_{f=f_{c0}}$ $=\frac{d}{df}\log|G(j2\pi f)|\Big{|}_{f=f_{c0}}$
$\quad+\frac{d}{df}\log|C(j2\pi f)|\Big{|}_{f=f_{c0}}$
@@
Under typical assumptions where $C(s)$ has integrator/low-pass behavior and $C_{\operatorname{delay}}(s)$ provides lead compensation:
$\left|\frac{d}{df}\log|L(j2\pi f)|\Big{|}_{f=f_{c0}}\right|\approx\gamma>0$ (183)
@@
The nominal phase margin without delay: $\phi_{\operatorname{design}}=180\textdegree+\arg L_{0}(j2\pi f_{c0})$
@@
Typical values: $\sigma=10\textdegree-15\textdegree$ for robust performance. $\square$
Appendix C High-Dimensional Topological Extensions
C-A Multi-Dimensional PH Stability (Complete Proof)
@@
$d_{\operatorname{PH}}^{(0:3)}(G_{\mathcal{C}}(t),G_{\mathcal{C}}(t+\delta))\leq\sum_{k=0}^{3}\alpha_{k}\left(C_{1,k}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+C_{2,k}\eta(\mathcal{L}_{\operatorname{ph}})\right)$ (191)
@@
$\|f_{t}-f_{t+\delta}\|_{\infty}$ $\leq L_{c}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+\eta(\mathcal{L}_{\operatorname{ph}})$ (197)
@@ Conclusion:
$d_{\operatorname{PH}}^{(0:3)}\leq\sum_{k=0}^{3}\alpha_{k}\left(C_{1,k}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+C_{2,k}\eta(\mathcal{L}_{\operatorname{ph}})\right)$ (198)
@@
C-B Multi-Scale Topological Stability (Complete Proof)
@@
$\leq L_{c}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+\eta(\mathcal{L}_{\operatorname{ph}})$ (206)
@@
C-C Probabilistic PH Stability (Complete Proof)
@@
Appendix D Complete Proofs of Main Theoretical Results
D-A ORTSF Continuity Proposition (Complete Proof)
@@
D-B Relational Consistency Theorem (Complete Proof)
@@
$\mathcal{L}_{\operatorname{context}}$ $=\mathcal{L}_{\operatorname{ricci\text{-}internal}}+\lambda_{\operatorname{ph}}\mathcal{L}_{\operatorname{ph}}$ (231)
@@
$\quad+C_{2,k}\eta(\mathcal{L}_{\operatorname{ph}})\Big{)}$ (234)
@@
$\mathcal{L}_{\operatorname{ph}}$ $<\frac{\eta(\epsilon)}{\lambda_{4}\lambda_{\operatorname{ph}}}$ (236)
@@
$\eta(\epsilon)$ $=\frac{\epsilon^{2}}{\lambda_{4}\left(\sum_{k=0}^{3}\alpha_{k}C_{1,k}\kappa\right)^{2}+\lambda_{4}\lambda_{\operatorname{ph}}\left(\sum_{k=0}^{3}\alpha_{k}C_{2,k}\right)^{2}}$ (237)
@@
$\quad\quad+\sum_{k=0}^{3}\alpha_{k}C_{2,k}\eta\left(\frac{\eta(\epsilon)}{\lambda_{4}\lambda_{\operatorname{ph}}}\right)$ (240)
@@
D-C BIBO Stability Theorem (Complete Proof)
@@
D-D Convergence Rate Theorem (Complete Proof)
@@
$=\mathcal{L}_{\operatorname{ricci\text{-}internal}}+\lambda_{\operatorname{ph}}\mathcal{L}_{\operatorname{ph}}$ (263)
@@
$\quad+C_{2,k}\eta(\mathcal{L}_{\operatorname{ph}}(k))\Big{)}$ (271)
@@
Since both $\mathcal{L}_{\operatorname{ricci\text{-}internal}}(k)$ and $\mathcal{L}_{\operatorname{ph}}(k)$ converge at rate $O(k^{-1/2})$:
@@
D-E Enhanced Unified Stability Bound (Complete Proof)
@@
$\quad+L_{\operatorname{ORTSF}}\eta(\mathcal{L}_{\operatorname{ph}})$ (280)
@@
$\quad+C_{2,k}\eta(\mathcal{L}_{\operatorname{ph}})\Big{)}$ (283)
@@
$\quad\leq L_{c}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+\eta(\mathcal{L}_{\operatorname{ph}})$ (285)
@@
$\|\mathcal{R}_{\operatorname{trace}}(t)-\mathcal{R}_{\operatorname{trace}}(t-\Delta t)\|\leq\eta(\mathcal{L}_{\operatorname{ph}})$ (290)
@@
$\quad+\sum_{k=0}^{3}\alpha_{k}C_{2,k}\eta(\mathcal{L}_{\operatorname{ph}})$ (296)
@@
$\quad+\eta(\mathcal{L}_{\operatorname{ph}})+L_{\operatorname{ORTSF}}\eta(\mathcal{L}_{\operatorname{ph}})$ (298)
@@
$\quad+\left(\sum_{k=0}^{3}\alpha_{k}C_{2,k}+1+L_{\operatorname{ORTSF}}\right)\eta(\mathcal{L}_{\operatorname{ph}})$ (301)
@@
$\quad+L_{\operatorname{ORTSF}}\eta(\mathcal{L}_{\operatorname{ph}})$ (304)

### v3

@@ Conclusion:
$d_{\operatorname{PH}}(G_{\mathcal{C}}(t),G_{\mathcal{C}}(t+\delta))\leq L_{c}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+C_{sem}\sqrt{\mathcal{L}_{\operatorname{context}}}$ (155)
@@
Patch-B: Discrete Predictor Continuity and Grönwall Bound (Complete Proof)
@@
$\|\mathcal{R}_{\operatorname{trace}}(t)-\mathcal{R}_{\operatorname{trace}}(t-\Delta t)\|$
$\qquad\leq C_{PH}\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+\eta(\mathcal{L}_{\operatorname{context}})$ (169)
@@
$\leq L_{\operatorname{total}}(C_{PH}\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+\eta(\mathcal{L}_{\operatorname{context}}))$ (171)
@@
Patch-D: Robust Phase Margin with Frequency Drift (Complete Proof)
@@
$\frac{d}{df}\log|L(j2\pi f)|\Big{|}_{f=f_{c0}}$ $=\frac{d}{df}\log|G(j2\pi f)|\Big{|}_{f=f_{c0}}$ (177)
$\quad+\frac{d}{df}\log|C(j2\pi f)|\Big{|}_{f=f_{c0}}$ (178)
@@
For model perturbation $\Delta G$:
$\Delta f_{c}\approx-\frac{\log|1+\Delta G(j2\pi f_{c0})/G(j2\pi f_{c0})|}{\frac{d}{df}\log|L(j2\pi f)|\Big{|}_{f=f_{c0}}}$ (180)
Under the assumption $\|\Delta G\|\ll\|G\|$:
$|\Delta f_{c}|\leq\frac{\|\Delta G\|/\|G\|}{|\frac{d}{df}\log|L(j2\pi f)||_{f=f_{c0}}|}\leq\alpha\|\Delta G\|$ (181)
for appropriately defined $\alpha>0$.
Step 4: Phase Margin Calculation
The effective phase margin becomes:
$\phi_{\operatorname{margin}}^{\operatorname{effective}}$ $=180^{\circ}+\arg[L(j2\pi f_{c})]$ (182)
$=180^{\circ}+\arg[G(j2\pi f_{c})]+\arg[C(j2\pi f_{c})]$ (183)
$\quad+\arg[C_{\operatorname{delay}}(j2\pi f_{c})]$ (184)
$=\phi_{\operatorname{design}}+\phi_{\operatorname{comp}}-360(f_{c}+\Delta f_{c})\Delta t-\epsilon$ (185)
where $\epsilon\geq 0$ accounts for higher-order uncertainty terms and unmodeled dynamics.
Step 5: Robustness Verification
For stability, we require:
Substituting the expression:
$\phi_{\operatorname{design}}+\phi_{\operatorname{comp}}-360(f_{c}+\Delta f_{c})\Delta t-\epsilon$ (187)
$\geq\phi_{\operatorname{safe}}+\sigma$ (188)
This gives the design constraint:
$\phi_{\operatorname{comp}}\geq 360(f_{c}+|\Delta f_{c}|)\Delta t+\epsilon+\phi_{\operatorname{safe}}+\sigma-\phi_{\operatorname{design}}$ (189)
Therefore, robust stability is achieved when the compensator phase contribution satisfies this bound. $\square$
Gauge Theory for Connection Consistency
Definition .1 (Gauge Group Action).
The gauge group $\mathcal{G}=\{g:V\to GL(d,\mathbb{R})\}$ acts on node variables by:
$(g\cdot f)_{i}=g_{i}f_{i},\quad\forall i\in V$ (190)
The connection transforms as:
$T_{ij}^{(g)}=g_{i}T_{ij}g_{j}^{-1}$ (191)
Theorem .2 (Gauge Invariance).
The connection Laplacian energy $\Phi(f)=\frac{1}{2}\sum_{(i,j)}w_{ij}\|T_{ij}f_{j}-f_{i}\|^{2}$ is gauge-invariant:
$\Phi(g\cdot f)=\Phi(f),\quad\forall g\in\mathcal{G}$ (192)
Proof.
Direct calculation:
$\Phi(g\cdot f)$ $=\frac{1}{2}\sum_{(i,j)}w_{ij}\|T_{ij}^{(g)}(g\cdot f)_{j}-(g\cdot f)_{i}\|^{2}$ (193)
$=\frac{1}{2}\sum_{(i,j)}w_{ij}\|g_{i}T_{ij}g_{j}^{-1}g_{j}f_{j}-g_{i}f_{i}\|^{2}$ (194)
$=\frac{1}{2}\sum_{(i,j)}w_{ij}\|g_{i}(T_{ij}f_{j}-f_{i})\|^{2}$ (195)
$=\frac{1}{2}\sum_{(i,j)}w_{ij}\|g_{i}\|^{2}\|T_{ij}f_{j}-f_{i}\|^{2}=\Phi(f)$ (196)
where the last equality uses $\|g_{i}\|=1$ for orthogonal gauge transformations. $\square$
Dynamic Topology and Nerve Complex Theory
Definition .3 (Nerve Complex Construction).
Given a cover $\mathcal{U}=\{U_{\alpha}\}$ of the configuration space, the nerve complex $\mathcal{N}(\mathcal{U})$ is the abstract simplicial complex with:
0-simplices: $\quad\{U_{\alpha}\mid\alpha\in I\}$ (197)
k-simplices: $\quad\{\{U_{\alpha_{0}},\ldots,U_{\alpha_{k}}\}\mid\bigcap_{i=0}^{k}U_{\alpha_{i}}\neq\emptyset\}$ (198)
Theorem .4 (Nerve Theorem for Semantic Covers).
If $\mathcal{U}$ is a good cover of the semantic configuration space (each $U_{\alpha}$ and all finite intersections are contractible), then:
$H_{*}(\bigcup_{\alpha}U_{\alpha})\cong H_{*}(\mathcal{N}(\mathcal{U}))$ (199)
This enables topological reasoning about semantic relationships through discrete simplicial structures.
Advanced Operator Theory for ONN
Definition .5 (Sectorial Operator).
A densely defined linear operator $A:D(A)\subset\mathcal{H}\to\mathcal{H}$ is sectorial with angle $\theta\in[0,\pi/2)$ if:
$\sigma(A)\subset S_{\theta}:=\{z\in\mathbb{C}\setminus\{0\}:|\arg(z)|\leq\theta\}$ (200)
and for all $\phi\in(\theta,\pi)$:
$\|(zI-A)^{-1}\|\leq\frac{C}{|z|},\quad z\notin S_{\phi}$ (201)
Theorem .6 (Fractional Powers and Interpolation).
For sectorial $A$ with $0\in\rho(A)$, fractional powers $A^{\alpha}$ for $\alpha>0$ are well-defined. The interpolation spaces satisfy:
$D(A),\mathcal{H}]_{\theta}=D(A^{\theta}),\quad 0<\theta<1$ (202)
This provides the analytical framework for smooth semantic transitions.
Stochastic Differential Equations on Manifolds
Definition .7 (Stratonovich SDE on Scene Graphs).
The semantic evolution follows the Stratonovich SDE:
$d\mathcal{S}_{t}=\nabla\circ dW_{t}+b(\mathcal{S}_{t},t)dt$ (203)
where $\nabla$ is the connection, $W_{t}$ is manifold-valued Brownian motion, and $\circ$ denotes Stratonovich integration.
Theorem .8 (Geometric Integration Preserves Structure).
Stratonovich integration preserves the geometric structure of the scene graph:
  1. 1.
    Metric preservation: $d\langle\mathcal{S},\mathcal{S}\rangle=2\langle d\mathcal{S},\mathcal{S}\rangle$
  2. 2.
    Connection preservation: $\nabla d\mathcal{S}=d(\nabla\mathcal{S})+R(dW,\mathcal{S})$
  3. 3.
    Curvature coupling: Noise interacts with curvature through the Ricci tensor
Homological Algebra for Constraint Systems
Definition .9 (Complex of Constraints).
The constraint system forms a cochain complex:
$0\to\mathcal{C}^{0}\xrightarrow{d_{0}}\mathcal{C}^{1}\xrightarrow{d_{1}}\mathcal{C}^{2}\to\cdots$ (204)
where $\mathcal{C}^{k}$ contains k-ary constraints and $d_{k}$ are constraint boundary maps.
Theorem .10 (Constraint Cohomology).
The cohomology groups $H^{k}(\mathcal{C}^{*},d)$ classify:
$H^{0}$ $=\text{Global consistency conditions}$ (205)
$H^{1}$ $=\text{Redundant constraint relations}$ (206)
$H^{2}$ $=\text{Higher-order constraint interactions}$ (207)
Non-trivial cohomology indicates fundamental constraint conflicts.
Spectral Graph Theory Extensions
Theorem .11 (Spectral Stability Under Perturbations).
For graph Laplacian $L$ with eigenvalues $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$, edge perturbations $\Delta E$ with $\|\Delta E\|\leq\varepsilon$ satisfy:
[…]
This ensures that semantic graph topology changes smoothly with small environmental perturbations.
Definition .12 (Effective Resistance Distance).
The effective resistance between vertices $i$ and $j$ is:
$R_{ij}=(e_{i}-e_{j})^{T}L^{+}(e_{i}-e_{j})$ (209)
where $L^{+}$ is the Moore-Penrose pseudoinverse of the graph Laplacian.
Theorem .13 (Resistance-Semantic Correspondence).
High effective resistance $R_{ij}$ correlates with semantic distance in the ONN embedding:
$\|\mathcal{S}_{i}-\mathcal{S}_{j}\|^{2}\geq\alpha R_{ij}-\beta$ (210)
for constants $\alpha,\beta>0$ depending on the connection weights and embedding dimension.
Information-Theoretic Bounds
Theorem .14 (Mutual Information and Topological Complexity).
The mutual information between scene components satisfies:
$I(\mathcal{S}_{i};\mathcal{S}_{j})\leq\log(\text{rank}(L_{1}))+H(\xi)$ (211)
where $H(\xi)$ is the entropy of the noise process and $\text{rank}(L_{1})$ captures topological complexity.
Definition .15 (Topological Entropy).
The topological entropy of the semantic dynamics is:
$h_{\text{top}}=\lim_{T\to\infty}\frac{1}{T}\log N_{T}(\varepsilon)$ (212)
where $N_{T}(\varepsilon)$ counts $\varepsilon$-separated orbits of length $T$.
Completion of Mathematical Framework
This appendix has established comprehensive mathematical foundations for all theoretical components referenced in the main text:
  • Topological Methods: Neck surgery, nerve complexes, persistent homology with rigorous distance bounds
  • Analytical Framework: Averaged operators, sectorial theory, fractional powers for smooth transitions
  • Geometric Structures: Gauge theory, connection consistency, Forman-Ricci curvature bounds
  • Stochastic Analysis: Stratonovich SDEs on manifolds preserving geometric structure
  • Control Theory: Complete phase margin analysis with frequency drift and model uncertainties
  • Homological Methods: Constraint complexes and cohomological classification of conflicts
  • Spectral Theory: Graph eigenvalue perturbation bounds and effective resistance metrics
  • Information Theory: Entropy bounds linking topology to information-theoretic complexity
These results provide the rigorous mathematical foundation supporting the seven-theorem framework and ensure that all referenced theoretical constructs have complete formal definitions and proofs.
@@
The nominal phase margin without delay: $\phi_{\operatorname{design}}=180^{\circ}+\arg L_{0}(j2\pi f_{c0})$
@@
Typical values: $\sigma=10^{\circ}-15^{\circ}$ for robust performance. $\square$
High-Dimensional Topological Extensions
Multi-Dimensional PH Stability (Complete Proof)
@@
$d_{\operatorname{PH}}^{(0:3)}(G_{\mathcal{C}}(t),G_{\mathcal{C}}(t+\delta))$ $\leq\sum_{k=0}^{3}\alpha_{k}\Big{(}C_{1,k}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}$
$\qquad+C_{2,k}\eta(\mathcal{L}_{\operatorname{context}})\Big{)}$ (220)
@@
$\|f_{t}-f_{t+\delta}\|_{\infty}$ $\leq L_{c}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+\eta(\mathcal{L}_{\operatorname{context}})$ (226)
@@
$d_{\operatorname{PH}}^{(0:3)}\leq\sum_{k=0}^{3}\alpha_{k}\left(C_{1,k}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+C_{2,k}\eta(\mathcal{L}_{\operatorname{context}})\right)$ (227)
@@
Multi-Scale Topological Stability (Complete Proof)
@@
$\leq L_{c}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+\eta(\mathcal{L}_{\operatorname{context}})$ (235)
@@
Probabilistic PH Stability (Complete Proof)
@@
Complete Proofs of Main Theoretical Results
ORTSF Continuity Proposition (Complete Proof)
@@
Relational Consistency Theorem (Complete Proof)
@@
$\mathcal{L}_{\operatorname{context}}$ $=\mathcal{L}_{\operatorname{ricci\text{-}internal}}+\lambda_{\operatorname{context}}\mathcal{L}_{\operatorname{context}}$ (260)
@@
$\quad+C_{2,k}\eta(\mathcal{L}_{\operatorname{context}})\Big{)}$ (263)
@@
$\mathcal{L}_{\operatorname{context}}$ $<\frac{\eta(\epsilon)}{\lambda_{4}\lambda_{\operatorname{context}}}$ (265)
@@
$\eta(\epsilon)$ $=\frac{\epsilon^{2}}{\lambda_{4}\left(\sum_{k=0}^{3}\alpha_{k}C_{1,k}\kappa\right)^{2}+\lambda_{4}\lambda_{\operatorname{context}}\left(\sum_{k=0}^{3}\alpha_{k}C_{2,k}\right)^{2}}$ (266)
@@
$\quad\quad+\sum_{k=0}^{3}\alpha_{k}C_{2,k}\eta\left(\frac{\eta(\epsilon)}{\lambda_{4}\lambda_{\operatorname{context}}}\right)$ (269)
@@
BIBO Stability Theorem (Complete Proof)
@@
Convergence Rate Theorem (Complete Proof)
@@
$=\mathcal{L}_{\operatorname{ricci\text{-}internal}}+\lambda_{\operatorname{context}}\mathcal{L}_{\operatorname{context}}$ (292)
@@
$\quad+C_{2,k}\eta(\mathcal{L}_{\operatorname{context}}(k))\Big{)}$ (300)
@@
Since both $\mathcal{L}_{\operatorname{ricci\text{-}internal}}(k)$ and $\mathcal{L}_{\operatorname{context}}(k)$ converge at rate $O(k^{-1/2})$:
@@
Enhanced Unified Stability Bound (Complete Proof)
@@
$\quad+L_{\operatorname{ORTSF}}\eta(\mathcal{L}_{\operatorname{context}})$ (309)
@@
$\quad+C_{2,k}\eta(\mathcal{L}_{\operatorname{context}})\Big{)}$ (312)
@@
$\quad\leq L_{c}\kappa\sqrt{\mathcal{L}_{\operatorname{ricci\text{-}internal}}}+\eta(\mathcal{L}_{\operatorname{context}})$ (314)
@@
$\|\mathcal{R}_{\operatorname{trace}}(t)-\mathcal{R}_{\operatorname{trace}}(t-\Delta t)\|\leq\eta(\mathcal{L}_{\operatorname{context}})$ (319)
@@
$\quad+\sum_{k=0}^{3}\alpha_{k}C_{2,k}\eta(\mathcal{L}_{\operatorname{context}})$ (325)
@@
$\quad+\eta(\mathcal{L}_{\operatorname{context}})+L_{\operatorname{ORTSF}}\eta(\mathcal{L}_{\operatorname{context}})$ (327)
@@
$\quad+\left(\sum_{k=0}^{3}\alpha_{k}C_{2,k}+1+L_{\operatorname{ORTSF}}\right)\eta(\mathcal{L}_{\operatorname{context}})$ (330)
@@
$\quad+L_{\operatorname{ORTSF}}\eta(\mathcal{L}_{\operatorname{context}})$ (333)
@@
Deleted Content: Rigorous Mathematical Foundations
Topological Neck Surgery Algorithm (Complete Formalization)
Definition (Unstable Topological Features): A 1-cycle $\gamma$ in $G_{\mathcal{C}}(t)$ is unstable if:
$\operatorname{pers}(\gamma)=\operatorname{death}(\gamma)-\operatorname{birth}(\gamma)<\epsilon_{\operatorname{neck}}$ (335)
where $\epsilon_{\operatorname{neck}}>0$ is a persistence threshold.
Algorithm 5 Discrete Topological Neck Surgery
0: Scene graph $G_{\mathcal{C}}(t)=(V,E)$, threshold $\epsilon_{\operatorname{neck}}>0$
0: Surgically corrected graph $G^{\prime}_{\mathcal{C}}(t)$
1:Phase 1: Neck Detection
[…]
3:for each 1-cycle $\gamma$ in $G_{\mathcal{C}}(t)$ do
4:  Compute $\operatorname{pers}(\gamma)$ via persistent homology
5:  if $\operatorname{pers}(\gamma)<\epsilon_{\operatorname{neck}}$ then
6:   $\mathcal{U}\leftarrow\mathcal{U}\cup\{\gamma\}$
7:  end if
8:end for
9:Phase 2: Curvature-Based Validation
[…]
11:for each $\gamma\in\mathcal{U}$ do
12:  $E_{\gamma}\leftarrow\{e\in E:e\text{ participates in cycle }\gamma\}$
13:  $\operatorname{Ric}_{\operatorname{avg}}\leftarrow\frac{1}{|E_{\gamma}|}\sum_{e\in E_{\gamma}}|\operatorname{Ric}_{F}(e)|$
14:  $\overline{\operatorname{Ric}}_{F}\leftarrow\frac{1}{|E|}\sum_{e\in E}\operatorname{Ric}_{F}(e)$
15:  if $\operatorname{Ric}_{\operatorname{avg}}>\overline{\operatorname{Ric}}_{F}+2\sigma_{\operatorname{Ric}}$ then
16:   $\mathcal{V}\leftarrow\mathcal{V}\cup\{\gamma\}$ {High curvature confirms instability}
17:  end if
18:end for
19:Phase 3: Surgical Correction
20:for each $\gamma\in\mathcal{V}$ do
21:  Find minimum cut edges $E_{\text{cut}}\subset E_{\gamma}$
22:  Apply surgical modification: $G^{\prime}_{\mathcal{C}}=(V,E\setminus E_{\text{cut}}\cup E_{\text{new}})$
23:  Update edge weights to minimize curvature variance
24:end for
25:Return corrected graph $G^{\prime}_{\mathcal{C}}(t)$
Theoretical Guarantees:
Theorem .16 (Neck Surgery Convergence).
The neck surgery algorithm terminates in finite steps and satisfies:
  1. 1.
    Monotonic Improvement: $\mathcal{L}_{\operatorname{ricci\text{-}internal}}^{\text{new}}\leq\mathcal{L}_{\operatorname{ricci\text{-}internal}}^{\text{old}}$
  2. 2.
    Connectivity Preservation: $\lambda_{2}(\mathcal{L}^{\text{new}})\geq\lambda_{2}(\mathcal{L}^{\text{old}})$
  3. 3.
    Essential Feature Preservation: All cycles with $\operatorname{pers}(\gamma)\geq\epsilon_{\operatorname{neck}}$ are preserved
Multi-Scale Filtration Framework (Complete Theory)
Definition (Scale-Space Filtration): Let $\Sigma=\{\sigma_{1}<\sigma_{2}<\cdots<\sigma_{m}\}$ be a finite set of scale parameters. For each $\sigma\in\Sigma$, define the smoothed filtration function:
$f_{t}^{(\sigma)}(e_{ij})=\Phi_{\sigma}\big{(}f_{t}(e_{ij})\big{)}$ (336)
where $\Phi_{\sigma}:\mathbb{R}\to\mathbb{R}$ is a $\sigma$-parametrized smoothing operator.
Algorithm 6 Multi-Scale Topological Analysis
0: Graph $G(V,E)$, scale set $\Sigma=\{\sigma_{1},\ldots,\sigma_{m}\}$, weights $\{w_{\sigma}\}$
0: Multi-scale persistent homology $\{D_{k}^{(\sigma)}\}$ and stability bounds
1:Initialize filtration functions $f_{t}$ for all edges
2:for each scale $\sigma\in\Sigma$ do
3:  Step 1: Apply smoothing operator
4:  for each edge $e_{ij}\in E$ do
5:   $f_{t}^{(\sigma)}(e_{ij})\leftarrow\Phi_{\sigma}(f_{t}(e_{ij}))$
6:  end for
7:  Step 2: Compute persistent homology
8:  $D_{k}^{(\sigma)}\leftarrow\text{PersistentHomology}(f_{t}^{(\sigma)},k)$ for $k=0,1,2$
9:  Step 3: Calculate scale-specific losses
10:  $\mathcal{L}_{\text{ricci}}^{(\sigma)}\leftarrow\sum_{e\in E}(\operatorname{Ric}_{F}^{(\sigma)}(e))^{2}$
11:  $\mathcal{L}_{\text{context}}^{(\sigma)}\leftarrow\|\mathcal{C}^{(\sigma)}x^{(\sigma)}-\tau^{(\sigma)}\|^{2}$
12:end for
13:Step 4: Compute multi-scale loss
[…]
15:Step 5: Verify stability bound
16: Check: $\sup_{\sigma\in\Sigma}d_{B}(D_{k}^{(\sigma)}(t),D_{k}^{(\sigma)}(t+\delta))\leq\|f_{t}-f_{t+\delta}\|_{\infty}$
17:Return multi-scale analysis results
Properties of Smoothing Operators:
  1. 1.
    Identity at Scale Zero: $\Phi_{0}=\text{Id}$
  2. 2.
    Order Preservation: $f_{1}\leq f_{2}\Rightarrow\Phi_{\sigma}(f_{1})\leq\Phi_{\sigma}(f_{2})$
  3. 3.
    Lipschitz Continuity: $|\Phi_{\sigma}(x)-\Phi_{\sigma}(y)|\leq|x-y|$
  4. 4.
    Scale Monotonicity: $\sigma_{1}<\sigma_{2}\Rightarrow$ more smoothing at $\sigma_{2}$
Multi-Scale Loss Function:
$\mathcal{L}_{\operatorname{context}}^{\operatorname{MS}}=\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}w_{\sigma}\left(\mathcal{L}_{\operatorname{ricci}}^{(\sigma)}+\lambda_{\operatorname{context}}\mathcal{L}_{\operatorname{context}}^{(\sigma)}\right)$ (337)
where $w_{\sigma}>0$ are scale-dependent weights with $\sum_{\sigma}w_{\sigma}=1$.
Theorem .17 (Multi-Scale Topological Stability).
Under the multi-scale framework:
$\sup_{\sigma\in\Sigma}d_{B}\big{(}D(f_{t}^{(\sigma)}),D(f_{t+\delta}^{(\sigma)})\big{)}\leq\|f_{t}-f_{t+\delta}\|_{\infty}$ (338)
Proof: Each $\Phi_{\sigma}$ is 1-Lipschitz, so bottleneck stability applies uniformly across scales. $\square$
Probabilistic Topological Stability (Complete Theory)
Definition (Sub-Gaussian Filtration Perturbation): A random variable $X$ is sub-Gaussian with parameter $\sigma^{2}$ if for all $t\in\mathbb{R}$:
$\mathbb{E}\left[\exp\left(\frac{t(X-\mathbb{E}[X])}{\sigma}\right)\right]\leq\exp\left(\frac{t^{2}}{2}\right)$ (339)
Assumption (Stochastic Filtration Model): The filtration perturbation $\xi_{t}=\|f_{t}-f_{t+\delta}\|_{\infty}$ satisfies:
$\mathbb{E}\left[\exp\left(\frac{t(\xi_{t}-\mathbb{E}[\xi_{t}])}{\sigma_{f}}\right)\right]\leq\exp\left(\frac{t^{2}}{2}\right)$ (340)
for some variance parameter $\sigma_{f}^{2}>0$.
Theorem .18 (Probabilistic PH Stability with Concentration).
Under the sub-Gaussian assumption:
$\mathbb{P}\left(d_{\operatorname{PH}}(G_{\mathcal{C}}(t),G_{\mathcal{C}}(t+\delta))>\varepsilon\right)$ (341)
$\quad\leq 2\exp\left(-\frac{(\varepsilon-L_{c}\mathbb{E}[\xi_{t}])^{2}}{2L_{c}^{2}\sigma_{f}^{2}}\right)$ (342)
Complete Proof: Step 1: By bottleneck stability: $d_{\operatorname{PH}}\leq L_{c}\xi_{t}$ Step 2: Sub-Gaussian tail bound: $\mathbb{P}(\xi_{t}>\mathbb{E}[\xi_{t}]+u)\leq\exp(-u^{2}/(2\sigma_{f}^{2}))$ Step 3: Set $u=(\varepsilon-L_{c}\mathbb{E}[\xi_{t}])/L_{c}$ and apply the bound.
Confidence Intervals: For confidence level $1-\alpha$:
$\varepsilon_{1-\alpha}=L_{c}\mathbb{E}[\xi_{t}]+L_{c}\sigma_{f}\sqrt{2\ln(2/\alpha)}$ (343)